# WWAggr: A Window Wasserstein-based Aggregation for Ensemble Change Point Detection

### v1

@@ Abstract
Index Terms:
Change Point Detection, Deep Ensembles, Wasserstein Distance, Model Calibration
@@ I Introduction
Change Point Detection (CPD) addresses the challenge of precise identification of the moments when some statistical properties of data distribution undergo alterations. Such a problem emerges in various real-world scenarios: manufacturing process monitoring [1, 2], server logs [3], financial data analysis [4, 5], or video surveillance [6]. In such situations, an unexpected change often indicates an emergency that requires an immediate response. Thus, developing accurate CPD methods is essential to ensure the safe deployment of automatic systems in applications.
CPD possesses a rich theoretical foundation with optimal procedures proposed for both online and offline detection [7, 8]. However, these methods rely on strong assumptions about data distributions, input dimensionality, or sparsity of the change signal, limiting their applicability to real-world data [9]. In response, various deep CPD models have been proposed [1, 6, 10, 11, 12, 13, 14], targeting practical demands and offering data-driven solutions with fewer prior assumptions.
While state-of-the-art (SOTA) deep CPD models demonstrate near-perfect performance for simple, low-dimensional time series, they often struggle when dealing with complex, multivariate data, like video streams [6, 13]. This observation suggests that individual models may lack the expressive power required to capture the diverse range of change points (CPs) occurring in the real world. A potential solution to this problem is ensembling, which has proven to boost weak standalone predictors [15] and enhance model expressiveness [16]. Moreover, the inconsistency of ensemble base learners with each other provides natural uncertainty estimation [16, 17] and may serve as an additional source of information for identifying CPs. Despite current advancements, current CPD ensembling [5, 18, 19] still relies on simple base detectors unsuitable for high-dimensional data, ignores the task’s specifics, and disregards the issue of optimal alarm threshold selection.
Figure 1: Teaser of our new aggregation procedure for effective high-dimensional CPD. First, an ensemble of deep change point detectors predicts well-calibrated CP scores for each moment. Second, these scores are aggregated via WWAggr — our sliding-window Wasserstein procedure. Applied to the calibrated scores, WWAggr better reflects changes in ensemble predictions, yielding near-optimal performance for the pre-defined threshold of $0.5$
@@
Table I: Mean performance ranks for the considered aggregation procedures. The ranks are averaged across all the models and datasets under study. The best value is highlighted in bold.
Base model Aggregation Mean $F_{1}$ rank
Single - 4.8
[…]
@@
The entire pipeline implementation, including ensemble training, calibration, and WWAggr is available online [^1] https://anonymous.4open.science/r/wwagr-ensemble-cpd-E7FF.
@@ II Related work
Another key aspect of finalizing deep ensemble predictions is model calibration. Most researchers find the calibration enhances empirical performance [37, 38]. The logic behind this step is that diverse base learners may produce outputs with different distributions, thereby requiring some transformation before final decision-making. Additionally, [39] shows that proper model calibration increases the discrepancy between normal and abnormal data in anomaly detection. We hypothesize that it can be beneficial for CPD in a similar manner and, thus, suggest using post-hoc calibration [20], which applies deterministic transformations to model outputs. The key methods here include Platt Scaling [40], Temperature Scaling [41], and beta calibration [42].
@@ III Problem statement
$\mathcal{H}_{0}^{(t)}\colon f_{X_{t-\omega:t}}=f_{X_{t:t+\omega}};$ (1)
$\mathcal{H}_{1}^{(t)}\colon f_{X_{t-\omega:t}}\neq f_{X_{t:t+\omega}}.$
Here, $X_{t-\omega:t}=\{\mathbf{x}_{t^{\prime}}\}_{t^{\prime}=t-\omega}^{t-1}$ and $X_{t:t+\omega}=\{\mathbf{x}_{t^{\prime}}\}_{t^{\prime}=t}^{t+\omega-1}$ are two consecutive sub-series of size $\omega$ for some step $t$, and $f_{X}$ denotes the PDF for a sample $X$. Eventually, $\tau$ is the first moment when $\mathcal{H}_{0}^{(t)}$ is rejected at some confidence level.
@@ IV-B Ensemble prediction aggregation
[…]
where $\Gamma\left(\mathbb{P},\mathbb{Q}\right)$ is the set of joint probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mathbb{P}$ and $\mathbb{Q}$, and $p\in\mathbb{N}$ is the order parameter. While difficult to compute in the general case, this distance is much simpler when $d=p=1$.
[…]
with $F_{\mathbb{P}}$ and $F_{\mathbb{Q}}$ being the corresponding CDFs. Substituting true CDFs with the empirical ones, we can estimate the 1-Wasserstein distance between the given samples $X^{(n)}\sim\mathbb{P}$ and $Y^{(n)}\sim\mathbb{Q}$ of size $n$ in the following way:
@@
While the 1-Wasserstein distance could also be estimated for $d>1$, the results from [45] established that the convergence speed of this estimator to the true distance is $O\left(n^{-1/d}\right)$. Consequently, with the ensemble of $K$ models, a faster convergence can be achieved if we use one-dimensional ($d=1$) samples of size $n=\omega K$ rather than estimate the distance between two $K$-dimensional samples, each of size $\omega$. Moreover, employing a moderate window size $\omega$ is generally preferred, as it results in a smaller inherent detection delay. This observation motivates us to introduce the flattening step within Algorithm 1. The decision is also justified by the assumption that after calibration (discussed below), CP scores predicted by distinct models come from the same distribution.
@@
In our Algorithm 1, the samples $X^{(n)}$ and $Y^{(n)}$ are the flattened history and future windows $\mathbf{h}_{t}$ and $\mathbf{f}_{t}$. They consist of the predicted CP scores $P_{1:T}$, bounded to $[0,1]$ due to the nature of the considered base models (see Appendix B). Thus, for any $i$, $0\leq\left|X_{(i)}-Y_{(i)}\right|\leq 1$, and, consequently, $0\leq w_{t}=d(\mathbf{f}_{t},\mathbf{h}_{t})=\hat{W}_{1}\left(X^{(n)},Y^{(n)}\right)\leq 1$. Therefore, for WWAggr, meaningful possible thresholds belong to $[0,1]$, making our aggregation convenient in practice. Simultaneously, there are no such guarantees for MMD.
Practically, reducing sensitivity to hyperparameters, particularly $h$, is highly desirable. The ideal CPD procedure should work effectively enough for a fixed alarm threshold, e.g., $0.5$. Our findings indicate that for the WWAggr method, this capability can be achieved through proper model calibration.
@@
In this work, we resort to post-hoc calibration methods, which assume that the model of interest has already been trained and its parameters are frozen. Therefore, better calibration is achieved by post-processing the model’s outputs. Let $s$ denote the model’s logit for a particular input. The predicted probability $p$ is then commonly computed as the sigmoid function: $p(s)=\sigma(s)=(1+\exp(-s))^{-1}$. We settle on the beta calibration method [42], where the Beta distribution inspires the logit transformation:
@@ V Experiments
Our final pipeline follows Algorithm 1, with $d$ being the 1-Wasserstein distance and the beta-calibrated ensemble base learners $g_{\mathbf{w}_{k}}$. In this section, we evaluate our ensemble CPD approach in real-data scenarios with various underlying base detectors for both supervised and unsupervised scenarios.
Table II: Detection $F_{1}$-scores for the ensembles of supervised and unsupervised CPD models. The results are averaged by three runs and are given in the format $mean\pm std$. The best values are highlighted in bold; the second best are underlined.
Base model Aggregation Dataset
Yahoo Road Accidents Explosions
Supervised models
BCE single - 0.895 $\pm$ 0.038 0.336 $\pm$ 0.024 0.695 $\pm$ 0.058
[…]
Min 0.872 $\pm$ 0.043 0.337 $\pm$ 0.031 0.679 $\pm$ 0.047
Max 0.901 $\pm$ 0.028 0.353 $\pm$ 0.031 0.735 $\pm$ 0.050
Median 0.908 $\pm$ 0.019 0.351 $\pm$ 0.009 0.709 $\pm$ 0.038
WWAggr 0.901 $\pm$ 0.027 0.383 $\pm$ 0.024 0.773 $\pm$ 0.011
InDiD single - 0.871 $\pm$ 0.023 0.319 $\pm$ 0.032 0.560 $\pm$ 0.070
[…]
Min 0.867 $\pm$ 0.036 0.302 $\pm$ 0.028 0.568 $\pm$ 0.087
Max 0.888 $\pm$ 0.023 0.317 $\pm$ 0.020 0.593 $\pm$ 0.027
Median 0.878 $\pm$ 0.023 0.337 $\pm$ 0.028 0.559 $\pm$ 0.029
WWAggr 0.882 $\pm$ 0.026 0.407 $\pm$ 0.007 0.621 $\pm$ 0.043
Unsupervised models
TS-CP2 single - 0.855 $\pm$ 0.034 0.359 $\pm$ 0.017 0.498 $\pm$ 0.080
[…]
Min 0.851 $\pm$ 0.018 0.354 $\pm$ 0.008 0.574 $\pm$ 0.019
Max 0.865 $\pm$ 0.005 0.364 $\pm$ 0.031 0.565 $\pm$ 0.047
Median 0.873 $\pm$ 0.004 0.378 $\pm$ 0.011 0.582 $\pm$ 0.014
WWAggr 0.891 $\pm$ 0.021 0.391 $\pm$ 0.022 0.618 $\pm$ 0.036
SN-TS2Vec single - 0.774 $\pm$ 0.033 0.361 $\pm$ 0.020 0.535 $\pm$ 0.053
[…]
Min 0.753 $\pm$ 0.008 0.354 $\pm$ 0.008 0.564 $\pm$ 0.061
Max 0.738 $\pm$ 0.025 0.364 $\pm$ 0.031 0.563 $\pm$ 0.031
Median 0.765 $\pm$ 0.047 0.378 $\pm$ 0.011 0.563 $\pm$ 0.017
WWAggr 0.785 $\pm$ 0.016 0.384 $\pm$ 0.039 0.564 $\pm$ 0.052
@@ V-B Datasets
Although our primary focus is on high-dimensional data CPD, we also evaluate our approaches on simpler sequences. For this, we selected the Yahoo! dataset [4], a widely recognized CPD benchmark. Its diverse collection of univariate series with changes in trends, seasonality, and noise regimes presents a sufficient challenge even for deep CP detectors [10, 11, 12]. As the most complicated datasets, we use high-dimensional video data: Explosions and Road Accidents, presented in [6, 46]. Both of them consider detecting sophisticated changes of varying intensities, which are complicated by the significant noisiness of the change signal and a wide diversity of normal behavior patterns. The main datasets’ characteristics are summarized in Table III, while the detailed description is postponed to Appendix A.
@@ V-C Main results
Following previous works [6, 10, 11, 12], CPD models are evaluated using the standard $F_{1}$ metrics, which is based on task-specific interpretation of the confusion matrix elements. Note that here, we are comparing not models themselves but various prediction aggregation techniques within each ensemble. As “naive” procedures, we consider point-wise “mean”, “min”, “max”, and “median” aggregations [19].
Table II presents the main results for the supervised and unsupervised CPD ensembles. The average performance ranks according to the $F_{1}$-score for all the considered aggregations are presented in Table I. Based on the results, we conclude that:
@@
    Ensembles of deep change point detectors consistently outperform standalone models. Confirming prior results [5, 19], even naive aggregation is sufficient for low-dimensional Yahoo. However, for higher-dimensional video data, task-agnostic procedures are suboptimal, while our WWAggr provides superior performance.
@@
    Our WWAggr technique is model-agnostic, working effectively with ensembles of various supervised and unsupervised deep CPD models. In almost all cases, it boosts performance, optimizing the detection $F_{1}$-score.
@@ V-E2 Mixed ensembles
Additionally, prove the versatility of the WWAggr procedure by experimenting with the mixed ensembles. In more detail, we construct two types of mixed ensembles: supervised (consisting of five BCE and five InDiD models) and unsupervised (with five TS-CP2 and five SN-TS2Vec base learners). The results are presented in Table V. As expected, we observe that the performance quality of the mixed ensembles lies between the corresponding homogeneous ensembles (Section V-C). However, it is more important to note that our WWAggr aggregation technique remains superior in this case as well, thereby proving its universality and robustness to the underlying ensemble structure.

### v2

@@ I Introduction
Change Point Detection (CPD) addresses the challenge of precise identification of the moments when some statistical data distribution properties undergo alterations. Such a problem emerges in various real-world scenarios: manufacturing process monitoring [1, 2], server logs [3], financial data analysis [4, 5], or video surveillance [6]. In these situations, an unexpected change often indicates an emergency that requires an immediate response. Thus, developing accurate CPD methods is essential to ensure the safe deployment of automatic systems in applications.
CPD possesses a rich theoretical foundation with optimal procedures proposed for both online and offline detection [7, 8]. However, these methods rely on strong assumptions about data distributions, input dimensionality, or sparsity of the change signal, which limits their applicability to real-world data [9]. In response, various deep CPD models have been proposed [1, 6, 10, 11, 12, 13, 14], targeting practical demands and offering data-driven solutions with fewer prior assumptions.
While state-of-the-art (SOTA) deep CPD models demonstrate near-perfect performance for simple, low-dimensional time series, they often struggle when dealing with complex, multivariate data, like video streams [6, 13]. This observation suggests that individual models may lack the expressive power required to capture the diverse range of change points (CPs) occurring in the real world. A potential solution to this problem is ensembling, which has proven to boost weak standalone predictors [15] and enhance model expressiveness [16]. Moreover, the inconsistency of ensemble base learners with each other provides natural uncertainty estimation [16, 17, 18] and may serve as an additional source of information for identifying CPs. Despite current advancements, recent CPD ensembling [5, 19, 20] still relies on simple base detectors unsuitable for high-dimensional data, ignores the task’s specifics, and disregards the issue of optimal alarm threshold selection.
Figure 1: Teaser of our new aggregation procedure for effective high-dimensional change point detection. First, an ensemble of deep change point (CP) detectors predicts well-calibrated CP scores for each moment. Second, these scores are aggregated via WWAggr — our sliding-window Wasserstein procedure. Applied to the calibrated scores, WWAggr better reflects changes in ensemble predictions, yielding near-optimal performance for the pre-defined threshold of $0.5$.
@@
Table I: Mean performance ranks for the considered aggregation procedures. The ranks are averaged across all the models and datasets under study. For standalone models with no aggregation of predictions, the results are marked as “s.m.” The best value is highlighted in bold.
Aggregation Mean $F_{1}$ rank
None (s.m.) 4.8
Mean 3.1
@@
The entire pipeline implementation, including ensemble training, calibration, and WWAggr is available online at https://github.com/stalex2902/wwagr-ensemble-cpd.
@@ II Related work
Another key aspect of finalizing deep ensemble predictions is model calibration. Most researchers find the calibration enhances empirical performance [18, 38, 39]. The logic behind this step is that diverse base learners may produce outputs with different distributions, thereby requiring some transformation before final decision-making. Additionally, [40] shows that proper model calibration increases the discrepancy between normal and abnormal data in anomaly detection. We hypothesize that it can be beneficial for CPD in a similar manner and, thus, suggest using post-hoc calibration [21], which applies deterministic transformations to model outputs. The key methods here include Platt Scaling [41], Temperature Scaling [42], and beta calibration [43].
@@ III Problem statement
$\mathcal{H}_{0}^{(t)}\colon f_{X_{t-\omega:t}}=f_{X_{t:t+\omega}};\quad\mathcal{H}_{1}^{(t)}\colon f_{X_{t-\omega:t}}\neq f_{X_{t:t+\omega}}.$ (1)
Here, $X_{t-\omega:t}=\{\mathbf{x}_{t^{\prime}}\}_{t^{\prime}=t-\omega}^{t-1}$ and $X_{t:t+\omega}=\{\mathbf{x}_{t^{\prime}}\}_{t^{\prime}=t}^{t+\omega-1}$ are two consecutive sub-series of size $\omega$ for some step $t$, and $f_{X}$ denotes the probability density function for a sample $X$. Eventually, $\tau$ is the first moment when $\mathcal{H}_{0}^{(t)}$ is rejected at some confidence level.
@@ IV-B Ensemble prediction aggregation
$W_{p}(\mathbb{P},\mathbb{Q})=\left(\inf_{\pi\in\Gamma\left(\mathbb{P},\mathbb{Q}\right)}{\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}{\|\mathbf{x}-\mathbf{y}\|^{p}d\pi}}\right)^{\frac{1}{p}},$ (2)
where $\Gamma\left(\mathbb{P},\mathbb{Q}\right)$ is the set of joint probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mathbb{P}$ and $\mathbb{Q}$, and $p\in\mathbb{N}$ is the order parameter. While being difficult to compute in the general case, this distance is much simpler when $d=p=1$:
$W_{1}(\mathbb{P},\mathbb{Q})=\int_{\mathbb{R}}{\left|F_{\mathbb{P}}(x)-F_{\mathbb{Q}}(x)\right|dx},$ (3)
with $F_{\mathbb{P}}$ and $F_{\mathbb{Q}}$ being the corresponding cumulative density functions (CDFs). Substituting true CDFs with the empirical ones, we can estimate the 1-Wasserstein distance between the given samples $X^{(n)}\sim\mathbb{P}$ and $Y^{(n)}\sim\mathbb{Q}$ of size $n$ in the following way:
@@
While the 1-Wasserstein distance could also be estimated for $d>1$, the results from [46] established that the convergence speed of this estimator to the true distance is $O\left(n^{-1/d}\right)$. Consequently, with the fixed ensemble of $K$ models, a faster convergence can be achieved if we use one-dimensional ($d=1$) samples of size $n=\omega K$ rather than estimate the distance between two $K$-dimensional samples, each of size $\omega$. Moreover, employing a moderate window size $\omega$ is generally preferred, as it results in a smaller inherent detection delay. This observation motivates us to introduce the flattening step within Algorithm 1. The decision is also justified by the assumption that after calibration (discussed below), CP scores predicted by distinct models come from the same distribution.
@@
In our Algorithm 1, the samples $X^{(n)}$ and $Y^{(n)}$ are the flattened history and future windows $\mathbf{h}_{t}$ and $\mathbf{f}_{t}$. They consist of the predicted CP scores $P_{1:T}$, bounded to $[0,1]$ due to the nature of the considered base models (see Appendix B). Thus, for any $i$, $0\leq\left|X_{(i)}-Y_{(i)}\right|\leq 1$, and, consequently, $0\leq w_{t}=d(\mathbf{f}_{t},\mathbf{h}_{t})=\hat{W}_{1}\left(X^{(n)},Y^{(n)}\right)\leq 1$. Therefore, for WWAggr, meaningful threshold values belong to $[0,1]$, making our aggregation convenient in practice. Simultaneously, there are no such guarantees for MMD.
Practically, reducing sensitivity to hyperparameters, particularly $h$, is highly desirable. The ideal CPD procedure should work effectively enough for a fixed alarm threshold, e.g., $0.5$. Our findings indicate that for the WWAggr method, this feature can be achieved through proper model calibration.
@@
To present a unified framework, independent of base model training procedures, we resort to post-hoc calibration methods, which assume that the model of interest has already been trained and its parameters are frozen. Therefore, better calibration is achieved by post-processing the model’s outputs. Let $s$ denote the model’s logit for a particular input. The predicted probability $p$ is then commonly computed as the sigmoid function: $p(s)=\sigma(s)=(1+\exp(-s))^{-1}$. We settle on the beta calibration method [43], where the Beta distribution inspires the form of the logit transformation:
@@ V Experiments
Our final pipeline follows Algorithm 1, with $d(\cdot,\cdot)$ being the 1-Wasserstein distance and base learners $g_{\mathbf{w}_{k}}$ being the beta-calibrated deep CPD models. In this section, we evaluate our approach in real-data scenarios with various underlying base detectors for both supervised and unsupervised scenarios.
Table II: Detection $F_{1}$-scores for the ensembles of supervised and unsupervised CPD models. The results are averaged by three runs and are given in the format $mean\pm std$. For standalone models, no prediction aggregation is used and these results are marked as “s.m.” The best values are highlighted in bold; the second best are underlined.
Dataset Aggregation Supervised models Unsupervised models
BCE InDiD TS-CP2 SN-TS2Vec
Yahoo None (s.m.) 0.895 $\pm$ 0.038 0.871 $\pm$ 0.023 0.855 $\pm$ 0.034 0.774 $\pm$ 0.033
Mean 0.892 $\pm$ 0.029 0.873 $\pm$ 0.025 0.872 $\pm$ 0.011 0.765 $\pm$ 0.022
Min 0.872 $\pm$ 0.043 0.867 $\pm$ 0.036 0.851 $\pm$ 0.018 0.753 $\pm$ 0.008
Max 0.901 $\pm$ 0.028 0.888 $\pm$ 0.023 0.865 $\pm$ 0.005 0.738 $\pm$ 0.025
Median 0.908 $\pm$ 0.019 0.878 $\pm$ 0.023 0.873 $\pm$ 0.004 0.765 $\pm$ 0.047
WWAggr 0.901 $\pm$ 0.027 0.882 $\pm$ 0.026 0.891 $\pm$ 0.021 0.785 $\pm$ 0.016
Explosions None (s.m.) 0.695 $\pm$ 0.058 0.560 $\pm$ 0.070 0.498 $\pm$ 0.080 0.535 $\pm$ 0.053
Mean 0.701 $\pm$ 0.052 0.588 $\pm$ 0.019 0.587 $\pm$ 0.044 0.563 $\pm$ 0.026
Min 0.679 $\pm$ 0.047 0.568 $\pm$ 0.087 0.574 $\pm$ 0.019 0.564 $\pm$ 0.061
Max 0.735 $\pm$ 0.050 0.593 $\pm$ 0.027 0.565 $\pm$ 0.047 0.563 $\pm$ 0.031
Median 0.709 $\pm$ 0.038 0.559 $\pm$ 0.029 0.582 $\pm$ 0.014 0.563 $\pm$ 0.031
WWAggr 0.773 $\pm$ 0.011 0.621 $\pm$ 0.043 0.618 $\pm$ 0.036 0.564 $\pm$ 0.052
Road Accidents None (s.m.) 0.336 $\pm$ 0.024 0.319 $\pm$ 0.032 0.359 $\pm$ 0.017 0.361 $\pm$ 0.020
Mean 0.354 $\pm$ 0.007 0.317 $\pm$ 0.023 0.381 $\pm$ 0.014 0.379 $\pm$ 0.009
Min 0.337 $\pm$ 0.031 0.302 $\pm$ 0.028 0.354 $\pm$ 0.008 0.354 $\pm$ 0.008
Max 0.353 $\pm$ 0.031 0.317 $\pm$ 0.020 0.364 $\pm$ 0.031 0.364 $\pm$ 0.031
Median 0.351 $\pm$ 0.009 0.337 $\pm$ 0.028 0.378 $\pm$ 0.011 0.378 $\pm$ 0.011
WWAggr 0.383 $\pm$ 0.024 0.407 $\pm$ 0.007 0.391 $\pm$ 0.022 0.384 $\pm$ 0.039
@@ V-B Datasets
Although our primary focus is on high-dimensional data CPD, we also evaluate our approaches on simpler sequences. For this, we selected the Yahoo! dataset [4], a widely recognized CPD benchmark. Its diverse collection of univariate series with changes in trends, seasonality, and noise regimes presents a sufficient challenge even for deep CP detectors [10, 11, 12]. As more complicated datasets, we use high-dimensional video data: Explosions and Road Accidents, presented in [6, 47]. Both of them consider detecting sophisticated changes of varying intensities, which are complicated by the significant noisiness of the change signal and a wide diversity of normal behavior patterns. The main datasets’ characteristics are summarized in Table III, while the detailed description is postponed to Appendix A.
@@ V-C Main results
Following previous works [6, 10, 11, 12], CPD models are evaluated using the standard $F_{1}$-score, which is based on task-specific interpretation of the confusion matrix elements. Note that here, we are comparing not models themselves but various prediction aggregation techniques within each ensemble. As “naive” procedures, we consider point-wise “mean”, “min”, “max”, and “median” aggregations [20].
Table II presents the main results for the supervised and unsupervised CPD ensembles. The average performance ranks for all the considered aggregations are presented in Table I. Based on the results, we conclude that:
@@
    Ensembles of deep change point detectors consistently outperform standalone models. Confirming prior results [5, 20], even simple aggregation is sufficient for low-dimensional Yahoo. However, for high-dimensional video data, these naive procedures are suboptimal, while our WWAggr provides superior performance.
@@
    Our WWAggr technique is model-agnostic, working effectively with ensembles of various supervised and unsupervised deep CPD models. In almost all cases, it boosts the performance.
@@ V-E2 Mixed ensembles
Additionally, we prove the versatility of the WWAggr procedure by experimenting with the mixed ensembles. In more detail, we construct two types of mixed ensembles: supervised (consisting of five BCE and five InDiD models) and unsupervised (with five TS-CP2 and five SN-TS2Vec base learners). The results are presented in Table V. As expected, we observe that the performance quality of the mixed ensembles lies between the corresponding homogeneous ensembles (Section V-C). However, it is more important to note that our WWAggr aggregation technique remains superior in this case as well, thereby proving its universality and robustness to the underlying ensemble structure.